What is the RMS speed of Helium atoms when the temperature of the Helium gas is 448 K?

(Possibly useful constants: the atomic mass of Helium is 4.00 AMU, the Atomic Mass Unit is: 1 AMU = 1.66×10-27 kg, Boltzmann's constant is: kB = 1.38×10-23 J/K.)

What would be the RMS speed, if the temperature of the Helium gas was doubled?
Answer #1

According to kinetic theory

Vrms = (3kB*T/m)1/2

={(3*1.38×10-23 J/K *448K) / (4*1.66×10-27 kg)}1/2

= 1671.303 m/s

if temp doubled

Vnewrms = (\sqrt{}2) * Vrms = 1.44*1671.303 = 2363.579 m/s